OpenStudy (anonymous):

What type of quadratic equation is represented in the graph below? Answer Perfect Square Trinomial Non-factorable Trinomial Not enough information Difference of Two Squares

OpenStudy (ash2326):

@AshleyF please post the graph

OpenStudy (ash2326):

@AshleyF tell me the two points where the curve is intersecting the x- axis?

OpenStudy (anonymous):

(0,-4)

OpenStudy (ash2326):

Isn't it intersecting at 2 and -2?

OpenStudy (anonymous):

o yeah my bad

OpenStudy (ash2326):

so what are the two roots?

OpenStudy (anonymous):

-2 and 2 so does it make it differences of two squares

OpenStudy (anonymous):

its a perfect square quadratic equation

OpenStudy (ash2326):

Yeah you're right great work:D

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:D

OpenStudy (ash2326):

@Siddharth18 for a perfect square, both roots need to be the same!!

OpenStudy (anonymous):

@ash2326,, sorry i was in a hurry to answer and so didn't notice it. now, this graph represents a quadratic eqn without any x term(i mean coefficient of x=0).so it reduces to the form ax^2-c=0 for some constants a and c with c/a>0. therefore x^2=c/a, this gives us two roots equal in magnitude but having opposite signs. by looking at the graph one can easily see that it represents x^2-4=0 giving the two roots as 2,-2

OpenStudy (anonymous):

@ash2326 thanks for pointing out the blunder